# Conventional and practical metallic superconductivity arising from repulsive Coulomb coupling

### v1

@@ II Dynamically screened Coulomb interaction
Consider a three-dimensional electron gas with dispersion $\varepsilon_{\mathbf{p}}=p^{2}/2m$ and spin degeneracy 2. At zero temperature, the Lindhard susceptibility [13]
@@
can be calculated analytically ($f_{\mathbf{p}}=\theta(\varepsilon_{\mathbf{p}}-\mu)$). Substitute $iq_{n}=\omega+i\epsilon,\epsilon\to 0$ for the retarded response and define dimensionless variables $\tilde{q}=q/2k_{F}$, $\tilde{\omega}=\omega/4\varepsilon_{F}$ and $\tilde{\chi}=\chi/d(\varepsilon_{F})$ (where $d(\varepsilon_{F})$ is the density of states at the Fermi energy $\varepsilon_{F}$, and $k_{F}$ is the corresponding Fermi momentum), we have [14, 15]
@@
Under the random phase approximation (RPA), the dielectric function is approximated by [15, 16, 17, 18]
@@
One may think that this might be an artifact of RPA, although RPA is the only available controlled approximation for the dielectric function, being a systematic expansion in $r_{s}$ summing up the most divergent ring diagrams in each order [18]. Here, we use the Hubbard local field correction to the RPA [19]. That is, instead of Eq. (5), the dielectric function is now approximated by [15]
@@
Since the large attractive regime in the energy-momentum space (the blue color in the real part plots) is directly extended from the plasmon pole (the line between blue and orange), a natural question is whether the plasmon pole approximation (PPA) itself is enough to reproduce the same qualitative result. The PPA is an approximation to the RPA where the response is represented entirely by a pole in the real part of response function (so the imaginary part becomes a delta function at this plasmon pole). The PPA is used extensively in the literature and is known to reproduce RPA results quantitatively [20, 21, 22]. Under the plasmon pole approximation, the dielectric function is [14]
@@
The results above clearly show that there is indeed a large regime in the energy-momentum space where the dynamically screened repulsive Coulomb interaction of the electrons produces an effective attractive interaction, making the discussion that whether it leads to a pairing instability relevant.
@@
Figure 4: Hydrodyanmic approximated Coulomb interaction $\tilde{u}_{\text{hydro}}(q,\omega)$, with dashed line indicate the damped region from RPA. As in 3, it reproduce the essential feature of Fig. 1 in the region to the left of the first dashed line.
@@ III BCS theory for plasmon-induced metallic superconductivity
The BCS theory is the paradigmatic theory for superconductivity with huge quantitative success for phonon-induced superconductivity in thousands of materials including all elemental metals (sometimes in the form of its strong-coupling extensions). Since the BCS theory is textbook material, we do not provide any redundant details, see, e.g., Refs. [17, 18, 15] (Details are provided in the next section, Sec. IV, where the rigorous Eliashberg-Migdal theory is discussed in the context of electron-electron interaction induced superconductivity.) In this section, we assume the applicability of the textbook BCS theory, and discuss its implications for $T_{c}$ in the plasmon induced superconductivity scenario of 3D jellium electron liquids.
The BCS theory provides the following formula for $T_{c}$ for a boson mediated SC, where the typical boson energy is $\omega_{b}$ and $\lambda_{b}$ is the dimensionless electron-boson coupling producing the SC:
@@
where we have neglected an unimportant constant of $O(1)$ in the prefactor. (We use units $\hbar=k_{B}=1$ throughout unless otherwise explicitly noted, so temperature/frequency/energy and momentum/wavenumber are the same in our notation.) For acoustic phonons, which are the primary boson-mediated mechanism driving SC in metals, the typical bosonic frequency $\omega_{b}\sim\omega_{D}$, where $\omega_{D}$ is the characteristic Debye energy for the phonons involved in the SC. The dimensionless electron-phonon coupling $\lambda_{b}$ in Eq. (13) varies between $0.1$ (Cu) to $1.7$ (Pb), and is defined by $\lambda_{b}=N(E_{F})V$, where $N$ and $V$ respectively denote the electronic density of states at the Fermi level and $V$ the electron-phonon coupling strength. All we need to do is to find the corresponding expressions for the parameters $\omega_{b}$ and $\lambda_{b}$ for the plasmons in the electron-plasmon superconductivity problem with the plasmons being the bosonic glue producing pairing.
@@
where $q_{p}=\omega_{p}/v_{F}$, with $v_{F}$ denoting Fermi velocity, is the cutoff momentum where the plasmon dispersion enters the electron-hole continuum (see Sec. II and Fig. 1), becoming damped. The plasmon no longer exists for $q>q_{p}$. Note that Eq. (19) is consistent with the estimates for the electron-plasmon coupling constant in other contexts [25].
@@
In Eq. (20), the $r_{s}^{-3/2}$ comes simply from the dependence of the plasmon energy on density: $\omega_{p}\sim n^{1/2}$ with $n\sim 1/r_{s}^{3}$ by definition. Eq. (20) with proper prefactors, which would be a model-dependent number, is the basic BCS prediction for plasmon induced $T_{c}$ in metals. The resultant $T_{c}$ (Fig. 5) has a peak around $r_{s}\sim 1$, and it vanishes quickly at high density (small $r_{s}$ because the basic coupling strength decreases as $r_{s}$ decreases producing an exponential drop in $T_{c}$) and slowly at low density (large $r_{s}$ because the plasmon frequency prefactor vanishes in a power law for large $r_{s}$ or low density as $n^{-1/2}$).
@@
One possibility is the well-known Coulomb repulsion effect (the $\mu^{*}$ effect”) [26]. The direct Coulomb repulsion between the electrons, e.g. the frequency independent exchange self-energy of Eq. (15) which depends only on the direct Coulomb repulsion, effectively reduces the attractive bosonic glue by an unknown amount since the full theory including the retarded attraction at the Fermi level with a strong static repulsion overall is intractable. But there are approximations leading to a phenomenological parameter $\mu^{*}$ which is considered to be $\mu^{*}\sim 0.1$$0.2$ for normal metals [26, 10]. This $\mu^{*}$ effectively suppresses the attractive glue, and crudely speaking, one could assume that the $\lambda_{b}$ is reduced by $\mu^{*}$ to produce an effective lambda, given by $\lambda^{*}\sim\lambda_{b}-\mu^{*}$. This, in fact, kind of explains why SC is absent in alkali metals such as Na and K or even Cu (where the phonon-induced $\lambda_{b}<0.2$, and hence the effective coupling is zero, producing no superconductivity).
@@
Could the $\mu^{*}$ effect save the situation for the plasmon-induced metallic superconductivity and provide a meaningful answer within the nominal BCS heuristics just by replacing $\lambda_{p}$ with $\lambda_{p}-\mu^{*}$. For $r_{s}\sim 6$, assuming $\mu^{*}\sim 0.2$, we have $\lambda_{p}\sim 1.2$ and $\lambda_{p}^{*}\sim 1.2-1=0.2$, still producing $T_{cp}\sim 10^{4}$ $\cdot\exp(-5)\sim 70 K$, still far too high (although much less ridiculous than $1000$ K). In addition, the basic theoretical conundrum is not resolved since for large enough $r_{s}$, the prediction would still be a high $T_{c}$, perhaps not $1000$ K, but $100$ K. For example, an interaction-driven 3D Wigner crystal happening at $r_{s}\sim 100$ would be a superconductor with $T_{cp}\sim 10$ K. In fact, such a crystalline superconductor would be a super-solid as it breaks both $U(1)$ and translational invariance.
@@
We believe that arbitrarily increasing $\mu^{*}$ by hand just in order to suppress plasmon-induced superconductivity at arbitrary $r_{s}$ is akin to saying that plasmons by itself cannot induce superconductivity in metals although the BCS theory (without any $\mu^{*}$ by definition) manifestly predicts a high-$T_{c}$ metallic superconductor. The real problem, as discussed in the next section, is that there is an additional conceptual theoretical issue that all current theories of superconductivity are based on the Midgal theorem, which asserts that vertex corrections are negligible thus ruling out all high frequency contributions to SC pairing. This theorem is valid for acoustic phonons where $\varepsilon_{F}\gg\omega_{D}$, but is simply invalid for plasmons since $\varepsilon_{F}\sim\omega_{p}$, and thus “integrating out of high frequency contribution” is neither allowed nor meaningful. This is because the superconductivity is arising from electrons themselves, thus fundamentally ruling out a separation between the low and high frequency contributions, which is the key to the Migdal theorem enabling an (in principle) quantitative theory for superconductivity, which is the Migfal-Eliashberg theory to be discussed in the next section.
@@
We conclude this section by asserting that although the naive BCS theory applied to the problem of the plasmon-induced metallic superconductivity predicts unreasonably high $T_{c}$, there is no resolution possible of this conundrum within the BCS-Migdal-Eliashberg paradigm since the Hamiltonian allows no separation of energy scales, and as such vertex corrections and $\mu^{*}$ effects are leading order effects and cannot be neglected as is done in all theories of superconductivity.

### v2

@@ II Dynamically screened Coulomb interaction
Consider a three-dimensional electron gas with dispersion $\varepsilon_{\mathbf{p}}=p^{2}/2m$ and spin degeneracy 2. At zero temperature, the Lindhard polarizability [Lindhard1954]
@@
can be calculated analytically ($f_{\mathbf{p}}=\theta(\varepsilon_{\mathbf{p}}-\mu)$). Substitute $iq_{n}=\omega+i\epsilon,\epsilon\to 0$ for the retarded response and define dimensionless variables $\tilde{q}=q/2k_{F}$, $\tilde{\omega}=\omega/4\varepsilon_{F}$ and $\tilde{\chi}=\chi/d(\varepsilon_{F})$ (where $d(\varepsilon_{F})$ is the density of states at the Fermi energy $\varepsilon_{F}$, and $k_{F}$ is the corresponding Fermi momentum), we have [Bruus2004, Mahan1980]
@@
Under the random phase approximation (RPA), the dielectric function is approximated by [Mahan1980, Hwang2018, schrieffer2018theory, Fetter2003]
@@
One may think that this might be an artifact of RPA, although RPA is the only available controlled approximation for the dielectric function, being a systematic expansion in $r_{s}$ summing up the most divergent ring diagrams in each order [Fetter2003]. Here, we use the Hubbard local field correction to the RPA [Hubbard1957]. That is, instead of Eq. (5), the dielectric function is now approximated by [Mahan1980]
@@
Figure 4: Hydrodyanmic approximated Coulomb interaction $\tilde{u}_{\text{hydro}}(q,\omega)$, with dashed line indicate the damped region from RPA. As in 3, it reproduces the essential feature of Fig. 1 in the region to the left of the first dashed line.
Since the large attractive regime in the energy-momentum space (the blue color in the real part plots) is directly extended from the plasmon pole (the line between blue and orange), a natural question is whether the plasmon pole approximation (PPA) itself is enough to reproduce the same qualitative result. The PPA is an approximation to the RPA where the response is represented entirely by a pole in the real part of response function (so the imaginary part becomes a delta function at this plasmon pole). The PPA is used extensively in the literature and is known to reproduce RPA results quantitatively [Lundqvist1967, Lundqvist1967a, Overhauser1971]. Under the plasmon pole approximation, the dielectric function is [Bruus2004]
@@
The results above clearly show that there is indeed a large regime in the energy-momentum space where the dynamically screened repulsive Coulomb interaction of the electrons produces an effective attractive interaction, making the discussion on that whether it leads to a pairing instability relevant.
Figure 5: 2D RPA Coulomb interaction $\tilde{u}^{\text{2D}}_{\text{RPA}}(q,\omega)$ with various $r_{s}$ and the top (bottom) row showing the real (imaginary) part. Dashed lines indicate the damped region. Note the different behavior of the plasmon dispersion from 3D.
We have also carried out similar calculations for the 2D electron gas in a jellium model (only for RPA and PPA), and the numerical results are shown in Figs.5 and 6 (for 2D RPA and PPA, respectively). We provide the 2D analytical results in the Appendix A. Again, RPA and PPA give similar results in the 2D case, demonstrating that the plasmon pole approximation captures the basic physics of dynamical screening very well. There is one glaring quantitative difference between 3D (Figs. 1 and 3) and 2D (Figs. 5 and 6). The attractive regions (in blue in all figures) in the momentum (wavevector)-energy (frequency) space for the screened interaction is drastically smaller in the 2D case than in the 3D case—the interaction is notably less attractive in 2D compared with 3D. This difference arises from the fact that the long wavelength, $q=0$, plasma frequency vanishes in 2D as $q^{1/2}$ since the 2D plasmon dispersion goes as $\omega_{p}(q)=(2\pi ne^{2}q/m)^{1/2}$ in the leading order (with $n$ as the 2D metallic electron density) [Stern1967], in contrast to the constant $q$-independent 3D plasma frequency. This considerably shrinks the regime between the plasmon branch $\omega_{p}(q)\sim q^{1/2}$ and the electron-hole continuum $\sim q$ before the plasmon hits the electron-hole continuum becoming damped. Since the screened interaction can only be negative or attractive for $\omega<\omega_{p}$, naturally the attractive region shrinks in 2D since the plasmon energy is bounded by the $q^{1/2}$ dispersion. Other than this quantitative difference, qualitatively 2D and 3D plots for the dynamically screened interactions are similar. One naive implication of this difference between 2d and 3D dynamically screened Coulomb interaction may, however, be relevant to the consideration of Coulomb interaction induced SC. This nominally implies that plasmon-induced SC is far less likely in 2D than in 3D since the effective electron-electron interaction is much less attractive (or more precisely, the phase space over which the interaction is attractive is much reduced) in 2D compared with 3D. In fact, a similar, but somewhat subtle, situation arises in the KL SC phenomena also, where naive arguments lead to a nonexistence of KL SC in 2D [Chubukov1993].
@@ III BCS theory for plasmon-induced metallic superconductivity
The BCS theory is the paradigmatic theory for superconductivity with huge quantitative success for phonon-induced superconductivity in thousands of materials including all elemental metals (sometimes in the form of its strong-coupling extensions). Since the BCS theory is textbook material, we do not provide any redundant details, see, e.g., Refs. [schrieffer2018theory, Fetter2003, Mahan1980] (Details are provided in the next section, Sec. IV, where the rigorous Eliashberg-Migdal theory is discussed in the context of electron-electron interaction induced superconductivity.) In this section, we assume the applicability of the textbook BCS theory, and discuss its implications for $T_{c}$ in the plasmon induced superconductivity scenario of 3D jellium electron liquids. Most, if not all, existing theories of the plasmon-induced SC in metals basically follow the simple BCS prescription. in some form or other.
The BCS theory provides the following formula for $T_{c}$ for boson mediated SC, where the typical boson energy is $\omega_{b}$ and $\lambda_{b}$ is the dimensionless electron-boson coupling producing the SC:
@@
where we have neglected an unimportant constant of $O(1)$ in the prefactor. (We use units $\hbar=k_{B}=1$ throughout unless otherwise explicitly noted, so temperature/frequency/energy and momentum/wavenumber/inverse wavelength are the same in our notation.) For acoustic phonons, which are the primary boson-mediated mechanism driving SC in metals, the typical bosonic frequency $\omega_{b}\sim\omega_{D}$, where $\omega_{D}$ is the characteristic Debye energy for the phonons involved in the SC. The dimensionless electron-phonon coupling $\lambda_{b}$ in Eq. (13) varies between $0.1$ (Cu) to $1.7$ (Pb), and is defined by $\lambda_{b}=N(E_{F})V$, where $N$ and $V$ respectively denote the electronic density of states at the Fermi level and $V$ the dimensionful electron-phonon coupling strength. All we need to do is to find the corresponding expressions for the parameters $\omega_{b}$ and $\lambda_{b}$ for the plasmons in the electron-plasmon superconductivity problem with the plasmons being the bosonic glue producing pairing.
Figure 6: 2D plasmon pole approximated Coulomb interaction $\tilde{u}^{\text{2D}}_{\text{PPA}}(q,\omega)$, with dashed line indicate the damped region from RPA. It reproduces the essential feature of Fig. 5
@@
where $q_{p}=\omega_{p}/v_{F}$, with $v_{F}$ denoting Fermi velocity, is the cutoff momentum where the plasmon dispersion enters the electron-hole continuum (see Sec. II and Fig. 1), becoming damped. The plasmon no longer exists for $q>q_{p}$. Note that Eq. (19) is consistent with the estimates for the electron-plasmon coupling constant in other contexts [Antropov1993].
@@
In Eq. (20), the $r_{s}^{-3/2}$ comes simply from the dependence of the plasmon energy on density: $\omega_{p}\sim n^{1/2}$ with $n\sim 1/r_{s}^{3}$ (in 3D) by definition. Eq. (20) with proper prefactors, which would be a model-dependent number, is the basic BCS prediction for plasmon induced $T_{c}$ in metals. The resultant $T_{c}$ (Fig. 7) has a peak around $r_{s}\sim 1$, and it vanishes quickly at high density (small $r_{s}$ because the basic coupling strength decreases as $r_{s}$ decreases producing an exponential drop in $T_{c}$) and slowly at low density (large $r_{s}$ because the plasmon frequency prefactor vanishes in a power law for large $r_{s}$ or low density as $n^{-1/2}$).
We have also carried out the BCS $T_{c}$ calculation in 2D following the same procedure as for the 3D case described above. The main differences between 2D and 3D are that in 2D the plasmon dispersion vanishes in the $q=0$ long wavelength limit and the 2D density of states and the static Thomas-Fermi screening wavevector $q_{\text{TF}}$ are constant independent of density: $\omega_{p}(q)=(2\pi ne^{2}q/m)^{1/2}$ in 2D in contrast to $\omega_{p}=(4\pi ne^{2}/m)^{1/2}$ in 3D, and $q_{\text{TF}}=2me^{2}/\hbar^{2}$ in 2D in contrast to $q_{\text{TF}}=(4me^{2}(3\pi^{2}n)^{1/3}/(\pi^{2}\hbar^{2}))^{1/2}$ in 3D, with $n$ being the 2D or 3D electron density as appropriate. These differences, however, do not affect the actual $T_{c}$ formula in the 2D case which is still given by the same equation as in the 3D case (i.e. Eq. (20) above) except that the dimensionless coupling parameter $r_{s}$ is now defined for the 2D case as $r_{s}=(\pi a_{0}^{2}n)^{-1/2}$, where $a_{0}=\hbar/e^{2}m$ is the Bohr radius and $n$ is the 2D electron density. Most of our discussions in this article are for 3D systems, but exactly the same considerations apply in the 2D case too.
For the sake of comparison with the corresponding phonon-induced SC in metals (2D or 3D), we provide in Fig. 8 the plots for the well-known BCS-Eliashberg formula for $T_{c}$ in metals as a function of the dimensionless electron-phonon coupling constant $\lambda_{\text{ph}}$ in units of the phonon frequency $\omega_{\text{ph}}$ assuming the accepted $\mu^{*}=0.15$ for the Eliashberg case [Allen1975]:
$\text{BCS: }T_{c}$ $\sim\omega_{\text{ph}}e^{-\frac{1}{\lambda_{\text{ph}}}}$ (21)
$\text{Eliashberg: }T_{c}$ $\sim\omega_{\text{ph}}e^{-\frac{1+\lambda_{\text{ph}}}{\lambda_{\text{ph}}-\mu^{*}(1+0.62\lambda_{\text{ph}})}}$ (22)
Note that for $\mu^{*}=0$ and $\lambda_{\text{ph}}\ll 1$, the BCS and Eliashberg formula coincide as they must since the BCS theory is a weak-coupling ($\lambda_{\text{ph}}\ll 1$) theory neglecting Coulomb repulsion ($\mu^{*}=0$). The main message from Figs. 7 and 8 are that the phonon-induced SC increases with the coupling eventually saturating at the effective phonon frequency for large coupling, but the plasmon-induced SC behaves qualitatively differently showing a maxima for an intermediate $r_{s}$ decreasing (exponentially) for $r_{s}\ll 1$ and in a power law manner for $r_{s}\gg 1$. This nonmonotonicity in the plasmon SC arises from the fact that the plasmons are inherent properties of the electrons themselves and not an independent bosonic mode as in the phonon case.
@@
Figure 8: The scaling of the $T_{c}$ of phonon-induced superconductivity with $\lambda_{\text{ph}}$ (unit not comparable with Fig. 7).
@@
One possibility is the well-known Coulomb repulsion effect (the $\mu^{*}$ effect”) [Allen1975], which has been already used in our Eq. (22) above for the phonon-mediated SC suppressed by Coulomb repulsion—this formula in Eq. (22) is the well-accepted Allen-Dynes empirical formula for metallic SC. The direct Coulomb repulsion between the electrons, e.g. the frequency independent exchange self-energy of Eq. (15) which depends only on the direct Coulomb repulsion, effectively reduces the attractive bosonic glue by an unknown amount since the full theory including the retarded attraction at the Fermi level with a strong static repulsion overall is intractable. But there are approximations leading to a phenomenological parameter $\mu^{*}$ which is considered to be $\mu^{*}\sim 0.1$$0.2$ for normal metals [Allen1975, Allen1983]. This $\mu^{*}$ effectively suppresses the attractive glue, and crudely speaking, one could assume that the $\lambda_{b}$ is reduced by $\mu^{*}$ to produce an effective lambda, given by $\lambda^{*}\sim\lambda_{b}-\mu^{*}$. This, in fact, kind of explains why SC is absent in alkali metals such as Na and K or even Cu (where the phonon-induced $\lambda_{b}<0.2$, and hence the effective coupling is zero, producing no superconductivity).
@@
Could the $\mu^{*}$ effect save the situation for the plasmon-induced metallic superconductivity and provide a meaningful answer within the nominal BCS heuristics just by replacing $\lambda_{p}$ with $\lambda_{p}-\mu^{*}$? For $r_{s}\sim 6$ (which is typical for a 3D metal), assuming $\mu^{*}\sim 0.2$, we have $\lambda_{p}\sim 1.2$ and $\lambda_{p}^{*}\sim 1.2-1=0.2$, producing $T_{cp}\sim 10^{4}$ $\cdot\exp(-5)\sim 70 K$, still far too high (although much less ridiculous than $1000$ K). In addition, the basic theoretical conundrum is not resolved since for large enough $r_{s}$, the prediction would still be a high $T_{c}$, perhaps not $1000$ K, but $100$ K. For example, an interaction-driven 3D Wigner crystal happening at $r_{s}\sim 100$ would be a superconductor with $T_{cp}\sim 10$ K. In fact, such a crystalline superconductor would be a super-solid as it breaks both $U(1)$ and translational invariance.
@@
We believe that arbitrarily increasing $\mu^{*}$ by hand just in order to suppress plasmon-induced superconductivity at arbitrary $r_{s}$ is akin to saying that plasmons by itself cannot induce superconductivity in metals although the BCS theory (without any $\mu^{*}$ by definition) manifestly predicts a high-$T_{c}$ metallic superconductor. The real problem, as discussed in the next section, is that there is an additional conceptual theoretical issue that all current theories of superconductivity are based on the Migdal theorem, which asserts that vertex corrections are negligible thus ruling out all high frequency contributions to SC pairing. This theorem is valid for acoustic phonons where $\varepsilon_{F}\gg\omega_{D}$, but is simply invalid for plasmons since $\varepsilon_{F}\sim\omega_{p}$, and thus “integrating out of high frequency contribution” is neither allowed nor meaningful. This is because the superconductivity is arising from electrons themselves, thus fundamentally ruling out a separation between the low and high frequency contributions, which is the key to the Migdal theorem enabling an (in principle) quantitative theory for superconductivity, which is the Migfal-Eliashberg theory to be discussed in the next section.
@@
We conclude this section by asserting that although the naive BCS theory applied to the problem of the plasmon-induced metallic superconductivity predicts unreasonably high $T_{c}$, there is no possible resolution of this conundrum within the BCS-Migdal-Eliashberg paradigm since the Hamiltonian allows no separation of energy scales, and as such vertex corrections and $\mu^{*}$ effects are leading order effects and cannot be neglected (or crudely approximated) as is done in all theories of superconductivity.